1.According to molecular orbital theory, the destructive interference of two atomic orbitals results in a(n)

2.According to molecular orbital theory, the constructive interference of two atomic orbitals results in a(n)
Answer #1

Answer:-

As we know that the atomic orbital of an atom can be expressed in terms of wave function ( )  i.e amplitude of an electron wave. In molecular orbital theory, the Linear combination of atomic orbitals (LCAO) method is used to explain the formation of molecules. when a molecule is formed, the atomic orbital (amplitude of an electron wave) of one atom (A) comes closer to other atomic orbital (amplitude of an electron wave) of the other atom (B), then in that condition the two electronic waves can be in phase (constructive interference) or out of phase (destructive interference). Thus there are two cases arise which are given below :-

Case-1

when two electronic waves of atomic orbital are in phase (constructive interference) so that they add up and the amplitude of the new wave is as follows:-

amplitude of the new wave () = A + B -----------------------------------------------(1)

where

A = amplitude of an electron wave of A atom i.e atomic orbital of A atom.

B = amplitude of an electron wave of B atom i.e atomic orbital of B atom.

Such type of interaction of two electronic waves of atomic orbital leads to the formation of Bonding molecular orbital in the molecular orbital theory. So equation (1) can be written as follows:-

amplitude of the new wave or Bonding molecular orbital  (BMO) = A + B

where

BMO = Bonding molecular orbital of the molecule.

A = amplitude of an electron wave of A atom i.e atomic orbital of A atom.

B = amplitude of an electron wave of B atom i.e atomic orbital of B atom.

Case-2

when two electronic waves of atomic orbital are in out of phase (destructive interference), the two waves try to cancel out the effect of each other so that the amplitude of the new wave is as follows:-

amplitude of the new wave () = A - B -----------------------------------------------(2)

where

A = amplitude of an electron wave of A atom i.e atomic orbital of A atom.

B = amplitude of an electron wave of B atom i.e atomic orbital of B atom.

Such type of interaction of two electronic waves of atomic orbital leads to the formation of Antibonding molecular orbital in the molecular orbital theory. So equation (2) can be written as follows:-

amplitude of the new wave or Antibonding molecular orbital  (ABMO) = A - B

where

ABMO = Antibonding molecular orbital of the molecule.

A = amplitude of an electron wave of A atom i.e atomic orbital of A atom.

B = amplitude of an electron wave of B atom i.e atomic orbital of B atom.

therefore

1.According to molecular orbital theory, the destructive interference of two atomic orbitals results in the formation of Antibonding molecular orbital (ABMO) .

2.According to molecular orbital theory, the constructive interference of two atomic orbitals results in the formation of Bonding molecular orbital (BMO) .

Note:- In molecular orbital theory, when two atomic orbitals overlaps then they forms one bonding molecular orbital and one antibonding molecular orbital.